OpenStudy (rea__):

Ernie deposited $4000 between two savings accounts. One account paid 4% simple interest and the other paid 2%. At the end of one year, Ernie had earned $140 in interest for both accounts. How much money did Ernie deposit in the account that paid 4% interest? A) $2000 B) $2500 C) $2800 D) $3000

OpenStudy (rea__):

@UliNK

OpenStudy (rea__):

$3000 x = amount at 4% 4000-x = amount at 2% .04(amt) + .02(amt) = 140 .04(x) + .02(4000-x) = 140 100[.04(x) + .02(4000-x) = 140] 4x + 2(4000-x) = 14000 4x + 8000 - 2x = 14000 2x = 6000 x = $3000 @ 4%

OpenStudy (rea__):

nvm i got it